Calculate the pressure exerted by a 0.25 mole sulfur hexafluoride in a steel vessel having a capacity of 1250 ml at 70.0°C

1
Expert's answer
2022-02-07T09:55:04-0500

Given:

"n=0.25\:\rm moles"

"V=\rm 1250\: ml=1.25*10^{-3}\: m^3"

"T=t+273=343\: \rm K"


The equation of state for ideal gas says

"p=\frac{nRT}{V}=\frac{0.25*8.31*343}{1.25*10^{-3}}=5.70*10^5\:\rm Pa"